OpenStudy (anonymous):

Simplify 5 square root 11 - 12 square root 11 - 2 square root 11

OpenStudy (anonymous):

\[5\sqrt{11} -12\sqrt{11} -2\sqrt{11}\]

OpenStudy (anonymous):

One second i will do it for you

OpenStudy (anonymous):

Your answer would come up too 742959540515

OpenStudy (anonymous):

I don't think i did it right because i did 5*5*5*5*5 and same thing for the others...